# Preference-free Alignment Learning with Regularized Relevance Reward

### v1

@@ Abstract
Learning from human preference has been considered key to aligning Large Language Models (LLMs) with human values. However, contrary to popular belief, our preliminary study reveals that reward models trained on human preference datasets tend to give higher scores to long off-topic responses than short on-topic ones. Motivated by this observation, we explore a preference-free approach utilizing ‘relevance’ as a key objective for alignment. On our first attempt, we find that the relevance score obtained by a retriever alone is vulnerable to reward hacking, i.e., overoptimizing to undesired shortcuts, when we utilize the score as a reward for reinforcement learning. To mitigate it, we integrate effective inductive biases into the vanilla relevance to regularize each other, resulting in a mixture of reward functions: Regularized Relevance Reward (R). R significantly improves performance on preference benchmarks by providing a robust reward signal. Notably, R does not require any human preference datasets (i.e., preference-free), outperforming open-source reward models in improving human preference. Our analysis demonstrates that R has advantages in elevating human preference while minimizing its side effects. Finally, we show the generalizability of R, consistently improving instruction-tuned models in various backbones and sizes without additional dataset cost 111Our code is available at github.com/naver-ai/RRR..
@@ 1 Introduction
To align large language models (LLM) with human values like helpfulness, human preference-based learning has been studied (Ouyang et al., 2022; Bai et al., 2022a, b; Rafailov et al., 2023). Typically, the preference is represented as a pair of responses where one is considered the chosen option, and the other is a rejected option for a given input under a Bradley-Terry model (Bradley & Terry, 1952). As illustrated in the upper side of Figure 1, Reinforcement Learning from Human Feedback (RLHF) is a dominant approach to exploiting pairwise feedback. The pairwise feedback is used to train a proxy reward model (RM), and a policy model is optimized over the reward signal from the RM using RL (Schulman et al., 2017; Ziegler et al., 2020). Also, there have been introduced alternative approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Sequence Likelihood Calibration (SLiC) (Zhao et al., 2023), which directly optimize preference pairs while not requiring an explicit RM.
@@
On the other hand, a line of research has studied the innate limitations of human preference (feedback). Xu et al. (2023) reveal prior knowledge affects the preference judgment in that crowd workers often choose the preferred response considering surface-level properties such as conciseness and specificity, while experts focus on more essential properties like factuality and completeness. Similarly, Hosking et al. (2023) further categorize the attributions of the preference judgments and confirm that human judgments could be biased by stylistic features like assertiveness. Similarly, during our preliminary study, we find that most open-source RMs (Köpf et al., 2023; Ethayarajh et al., 2023a; Touvron et al., 2023; Cui et al., 2023), trained with human preference, struggle with discriminating responses that are more related or on-topic with the given input, while an unsupervised retriever (Izacard et al., 2021) perfectly distinguishes them (Figure 2 shows an example.). This observation makes us wonder whether relevance can be used for the LLM alignment without explicit human preference, i.e., directly using it as a reward signal.
In this work, we present a preference-free approach focusing on relevance, motivated by our observation. We define ‘Relevance’ as the degree to which an assistant system’s responses are appropriate and directly connected to the specific needs or requests presented by the users. As shown in Figure 1, our approach does not require any preference dataset to align LLMs. Instead, we propose a mixture of reward functions: Regularized Relevance Reward (R). We show that when regularized appropriately, the relevance between the input and output could be a reliable learning signal. Specifically, we integrate a set of effective inductive biases such as length incentive, repetition penalty, and distinction by query type into the relevance score to suppress the risk of reward hacking, which is a phenomenon to exploit undesired shortcuts in the reward signal (Skalse et al., 2022; Pang et al., 2022). We use R as a reward function to align LLMs within Proximal Policy Optimization (PPO) (Schulman et al., 2017; Ziegler et al., 2020).
Experimental results show that our method significantly improves human preference over LLMs even without explicit human or AI feedback, outperforming state-of-the-art open-source RMs (Köpf et al., 2023; Ethayarajh et al., 2023a; Touvron et al., 2023; Cui et al., 2023). Also, we offer an in-depth analysis to compare the resulting response distribution between models trained with preference-based RM and our R, focusing on relevance. We find that all PPO models tend to generate irrelevant sentences as the response gets longer, but the model trained with R produces more relevant responses. Further ablation study demonstrates that R has advantages in improving preference while minimizing alignment tax (Ouyang et al., 2022), a phenomenon where the increased preference accompanies the degraded performance on other NLP tasks. Finally, R consistently improves aligned behaviors of strong instruction-following (SFT) models, Vicuna and Mistral (Chiang et al., 2023; Jiang et al., 2023; Tunstall et al., 2023), using only input-output pairs, i.e., preference-free.
2 Motivation and Preliminary Study
In alignment learning, human preference could be divided into several facets (Ye et al., 2023; Kim et al., 2023b), e.g., fluency, accuracy, completeness, relevance of the responses, etc. Among them, Relevance mainly refers to the degree to which the responses of an assistant system are appropriate and directly connected to the specific needs or requests presented by the users. It emphasizes the system’s ability to understand and appropriately address the user’s intent and context. Our study focuses on the response’s relevance with the provided query from the users, starting by asking the following two questions: (1) Can open-source RMs capture relevance? and (2) Can relevance-only be used for alignment?
2.1 Can open-source RMs capture relevance?
First, we study how open-source reward models (RM) capture relevance between query and response behind an overall preference. A common belief is that the reward model is capturing relevance well, but it turns out that (1) it does not, and (2) a simple relevance model like Contriever (Izacard et al., 2021) does it much better out of the box. For this study, we devise an adversarial test set, Synthetic Relevance (SynRel), containing 530 (query, relevant response, and irrelevant response) triplets, as illustrated in Figure 2. Specifically, we randomly sample 530 entities whose popularity is less than 1k in PopQA (Mallen et al., 2023) 222The PopQA dataset contains the quantified popularity of each entity, e.g., monthly page-view of Wikipedia.. The query for each instance is formed using the template “Please tell me about {entity}”, and the positive (chosen) response is created by compiling the ground truth property of the entity, e.g., “{entity} is {property}”. The negative (rejected) response is generated by prompting another negative entity to GPT-3.5-turbo (OpenAI, 2023).
We compare four RMs, OASST-RM 333OpenAssistant/reward-model-deberta-v3-large-v2 (Köpf et al., 2023), SteamSHP-XL (Ethayarajh et al., 2023a), LLaMA-2-7B-RM (Touvron et al., 2023), which is our implementation trained on Anthropic-HH (Bai et al., 2022a) and SHP (Ethayarajh et al., 2023a), and UltraRM-13B (Cui et al., 2023), on the SynRel test set. Also, we include Contriever (Izacard et al., 2021), originally used as a retriever, as a representative model trained to capture relevance. We report preference accuracy for selecting the chosen response for a given query in Table 1. We find that the test set of SynRel is challenging for most open-source RMs, even though they perform well in other preference datasets, Anthropic-HH and SHP. In other words, the reward models trained on overall human preference cannot capture the trivial relevance problems that a far smaller retriever model can perfectly distinguish 444The parameter size of UltraRM (13B) is about 118x larger than that of Contriever (110M)..
RM Overall Preference Relevance Anthropic-HH SHP SynRel Random 50.0 50.0 50.0 OASST-RM 67.6 53.9 52.6 SteamSHP-XL 68.5 74.8 58.9 LLaMA2-7B-RM 69.6 72.3 45.1 UltraRM-13B 70.8 73.4 59.8 Contriever 54.0 56.0 100
2.2 Can relevance-only be used for alignment?
From the observation of the previous subsection, we wonder if relevance could be used as a meaningful learning signal for LLM alignment. We could think of two ways to implement this. The first is training a canonical RM on relevance dataset (Wu et al., 2023), and the second is using an off-the-shelf retriever directly for the RM. Here, we choose the second option using Contriever denoted with for simplicity. Let be a query and be a rollout response in the exploration stage of Proximal Policy Optimization (PPO) (Schulman et al., 2017; Ziegler et al., 2020). We compute the relevance score between the query and sampled response: .
We use the Anthropic-HH (Bai et al., 2022a) for the experiments. We start by training supervised policy, i.e., supervised fine-tuning (SFT), using chosen responses in the dataset. Then, we perform PPO training upon the SFT model against relevance score . Please see Appendix C for more implementation details. Finally, We conduct GPT-4 evaluation (OpenAI, 2023) to inspect whether there are preference improvements after the RL training compared to the SFT model. Specifically, we compare the SFT and PPO models with a pair-wise judgment of Vicuna-Bench (Chiang et al., 2023). We report the adjusted win rate computed by (# Win + 0.5 # Tie) / # All) comparing the PPO models with the SFT baseline.
Model Vicuna-Bench % Adjusted Win vs. SFT Relevance score Avg # Tokens SFT 50.0 4.94 173 + PPO w. OASST-RM 61.3 6.49 174 w. UltraRM-13B 73.8 6.34 227 w. 20.0 8.06 69
In Table 2, we find that solely using can not improve the preference and even worsens the model. Contrary to other RMs, the relevance reward significantly drops the adjusted win rate in the Vicuna-Bench. We further analyze the problem by measuring query-response relevance score and response length. For this, we employ another open-source relevance model trained as a reranker 555huggingface.co/cross-encoder/ms-marco-MiniLM-L-12-v2 to compute the relevance score. The PPO model trained using the demonstrates the highest relevance score; however, its responses are notably brief in length. We speculate using only for the reward is vulnerable to reward hacking (Skalse et al., 2022; Pang et al., 2022). For example, the best strategy to maximize would be just copying the question for the response. In fact, about 40% of the generated responses are merely copies of the given questions. Also, it reminds us of the trade-off relationship between relevance and the short response length discussed in Wu et al. (2023).
3 R: Regularized Relevance Reward
To overcome the easily hackable nature of the vanilla relevance reward , we propose a Regularized Relevance Reward (R), the mixture of reward functions, including the that regularize each other. Note that this is just one way to regularize the vanilla relevance reward with a strong inductive bias, and one could explore better ways. We will demonstrate each function in this Section.
Length Incentive (LI)
@@ 3 R: Regularized Relevance Reward
First, we introduce Length Incentive (LI) to mitigate the ’s tendency to shorten the length of responses. It scores the reward by calculating . Intuitively, this reward function promotes the longer generation regardless of the given query . It is similar to LPPO suggested by concurrent work (Singhal et al., 2023). However, we do not normalize the reward with the target length. We find that combining LI with the is somewhat effective in improving preference. However, in this setting, we encounter another kind of reward hacking, undesired repetitions in the generations. For example, the best strategy to maximize the is just repeating the .
Repetition Penalty (RP)
@@
Hence, we additionally introduce the Repetition penalty (RP) to regularize the undesired outputs. It checks the unique trigram ratio of , i.e., . Integrating it with the reward functions efficiently prevents redundancy and encourages proper length of outputs. The intuition behind the mixture of reward functions is aiming for high relevance between query and generated response while promoting long response length and less repetitions. However, the Length Incentive could be problematic when the question requires constrained or factual responses. In these cases, consistency of response might be much more important than the long and diverse responses, as discussed in Touvron et al. (2023); Song et al. (2023b).
Reference Answer Relevance ()
@@
We apply the different relevance rewards according to query type to handle such cases. Specifically, we define two types of queries, Open-ended (OE) and Closed-ended (CE), according to whether they require creative and open-ended or consistent and constrained responses, similar to Song et al. (2023b). Then, we include reference answer relevance , where the is the ground-truth reference response of given query.
Finally, we present our final reward function design, Regularized Relevance Reward (R):
, where the indicates the pre-identified query types and the indicates a linear interpolation function mapping the range of to the range of for stable RL training. The mixture of reward functions regularizes each other, making it robust to reward hacking. Besides, It only requires an SFT dataset consisting of and pairs, eliminating the need for an extensive feedback dataset to train an additional reward model.
Query-Type Classification
@@
Deciding whether the given query requires open-ended generation may not be obvious because of its subjectivity in interpreting the range of ‘open-ended’. Nevertheless, we define the meaning of open-ended so that the responses to the query can be relatively anything and creative. For example, “How can I make a good first impression?” could be an open-ended query. On the other side, the closed-ended query indicates the corresponding responses should be consistent and constrained as a closed form. “How many movies are in the Fast and Furious franchise?” would be one of the closed-ended queries. Considering resource constraints, we automatically annotate the query type (OE or CE) used in R by prompting it to GPT-4 (OpenAI, 2023). We find that about 83% of decisions by GPT-4 are agreed upon when the authors manually validate 100 samples in a preliminary stage. Also, when we distill the classification task into in-house LLM, LLaMA-2-7B (Touvron et al., 2023), we achieve competitive accuracy in the test set 666When we do not specify otherwise, is derived from GPT-4 in our experiments.. Please see Appendix B for more details of the query-type classification.
PPO training against R
@@
We use the resulting mixture of reward functions, R, to further optimize an SFT model with reinforcement learning. Consequently, the following PPO objective induces the model to maximize R,
while online policy and the fixed reference (SFT) policy do not diverge too much by regularizing it with KL constraints (Schulman et al., 2017; Ziegler et al., 2020), where the is train dataset and the is the coefficient to control the KL penalty.
Model # Feedback to train RM Vicuna-Bench % Adjusted Win vs. SFT AlpacaEval % Win vs. Text-Davinci-003 MT-Bench 1st / 2nd / Avg ToxiGen () SFT (Anthropic-HH) - 50.0 15.9 4.6 / 3.4 / 3.9 2.6 + PPO w. SteamSHP-XL 477k 59.4 19.6 4.2 / 3.4 / 3.8 6.2 w. OASST 393k 61.3 19.1 4.7 / 3.1 / 3.9 5.4 w. LLaMA2-7B-RM 509k 62.5 19.8 4.6 / 3.5 / 4.0 5.6 w. UltraRM-13B 256k 73.8 28.0 4.9 / 3.6 / 4.2 4.7 w. R (Ours) 0 76.9 23.4 5.0 / 3.7 / 4.4 3.9

### v3

# Rethinking the Role of Proxy Rewards in Language Model Alignment

@@ Abstract
Learning from human feedback via proxy reward modeling has been studied to align Large Language Models (LLMs) with human values. However, achieving reliable training through that proxy reward model (RM) is not a trivial problem, and its behavior remained as a black-box. In this paper, we study the role of proxy rewards in the LLM alignment via ‘reverse reward engineering’ by composing interpretable features as a white-box reward function. We aim to replicate the ground truth (gold) reward signal by achieving a monotonic relationship between the proxy and gold reward signals after training the model using the proxy reward in reinforcement learning (RL). Our findings indicate that successfully emulating the gold reward requires generating responses that are relevant with enough length to open-ended questions, while also ensuring response consistency in closed-ended questions. Furthermore, resulting models optimizing our devised white-box reward show competitive performances with strong open-source RMs in alignment benchmarks. We highlight its potential usage as a simple but strong reward baseline for the LLM alignment, not requiring explicit pairwise human feedback dataset and RM training 111Our code is available at github.com/naver-ai/rethinking-proxy-reward..
@@ 1 Introduction
To align large language models (LLM) with human values like helpfulness, human feedback-based learning has been studied Ouyang et al. (2022); Bai et al. (2022a, b); Rafailov et al. (2023). Reinforcement Learning from Human Feedback (RLHF) is a dominant approach to exploiting human feedback. Typically, the human feedback is used to train a proxy reward model (RM), and a policy model is optimized over the reward signal from the RM using RL Schulman et al. (2017); Ziegler et al. (2020).
However, reliable RL training with the proxy RM is not a trivial problem. Gao et al. (2023) study the overoptimization problem of the policy training against the imperfect RM. It is a phenomenon that the proxy rewards from the RM are increased continuously, but the true rewards are saturated or even decreased in fact. Similarly, the policy model often finds undesired shortcuts in the reward signal from the imperfect black-box RM, i.e., “reward hacking” Skalse et al. (2022); Pang et al. (2022).
@@
A line of research has studied the innate limitations of human feedback. Xu et al. (2023) reveal prior knowledge affects the preference judgment in that crowd workers often choose the preferred response considering surface-level properties such as conciseness and specificity, while experts focus on more essential properties like factuality and completeness. Similarly, Hosking et al. (2023) further categorize the attributions of the preference judgments and confirm that human judgments could be biased by stylistic features like assertiveness. Moreover, Singhal et al. (2023) analyze the correlations between long response and RLHF, demonstrating a verbosity bias in the human feedback.
In this work, we show merely combining interpretable features as a proxy reward function can substantially maximize the score of state-of-the-art RM Lambert et al. (2024) and further enhance human preferences in various alignment benchmarks Chiang et al. (2023); Li et al. (2023); Zheng et al. (2023). Specifically, we study the role of proxy rewards learned by human feedback in the RLHF by employing “reverse reward engineering” with interpretable features, such as length or relevance of the responses, to compose the white-box reward function, as shown in Figure 1. We define the goal of reverse engineering as achieving a monotonic relationship between the proxy and ground truth (gold) reward signals after training the model with the proxy reward in an RL manner. This is to verify whether training to maximize the proxy white-box reward during RL training also maximizes the gold reward signal at the test time.
Experimental results indicate that the key to imitating the gold reward signal lies in generating responses that are sufficiently lengthy yet relevant and faithful for open-ended queries while ensuring consistency in responses to closed-ended queries. Contrary to previous study Singhal et al. (2023), our results demonstrate that solely optimizing towards lengthy response makes drastic drops in the gold reward, showing severe overoptimization. Also, the reward branching according to query type reliably improves the win rate while not introducing unnecessary verbosity compared to other baseline reward options. Furthermore, we find that the reward differentiation brings advantages in improving preference while minimizing alignment tax Ouyang et al. (2022), a phenomenon where the increased preference accompanies the degraded performance on other NLP tasks. Notably, the engineered white-box reward often results in even more improvements than the strong open-source RM like UltraRM-13B Cui et al. (2023) and also generally works well across LLM backbones, demonstrating its potential usage as a baseline reward.
Our contributions are summarized in three folds:
We investigate the role of proxy rewards learned by human feedback via reverse reward engineering with interpretable features.
Our results suggest that the key to replicating the gold reward involves producing responses that are not only detailed but also relevant for open-ended queries while maintaining consistency in responses to closed-ended queries.
We highlight the potential usage of the reverse-engineered white-box reward function as a simple but strong reward baseline not requiring a pairwise human feedback dataset.
2 Related Work
Reinforcement Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) is the most prevalent approach to leverage the human preference Ziegler et al. (2020); Ouyang et al. (2022); Bai et al. (2022a); Liu et al. (2022); Scheurer et al. (2023). These approaches include the reward modeling stage to develop a proxy RM predicting the human preference over the response, e.g., typically as a scalar format. Then, the trained proxy RM computes reward scores for the sampled responses in the RL stage, i.e., rollout. Recently, replacing human feedback with AI or synthetic feedback has been studied Bai et al. (2022b); Kim et al. (2023); Sun et al. (2023b); Cui et al. (2023). There are attempts to leverage multiple rewards beyond the overall reward Touvron et al. (2023); Wu et al. (2023); Rame et al. (2023); Jang et al. (2023). Wu et al. (2023) devise fine-grained rewards consisting of relevance, factuality, and information completeness to enhance long-form question-answering.
Pitfalls of Reward Modeling and Human Feedback
RLHF faces various challenges, especially in reward modeling and human feedback itself Casper et al. (2023). Gao et al. (2023) present the overoptimization problem by the imperfect proxy reward models. The issue arises when the policy, tailored to enhance the estimated rewards, fails to improve real-world rewards. Many studies try to mitigate the overoptimization problem Moskovitz et al. (2023); Coste et al. (2023). Similarly, Shen et al. (2023) study reward inconsistency and leverage a sentence encoder to increase the reward consistency. The work leverages SimCSE Gao et al. (2021) to augment the preference dataset and normalize reward scores. The policy models within the RLHF often find and exploit undesired shortcuts or errors in the RMs to maximize the proxy rewards. The phenomenon is known as reward hacking Amodei et al. (2016); Pan et al. (2022); Pang et al. (2022); Song et al. (2023); Eisenstein et al. (2023). Much literature points out that human preference can be easily biased by superficial features such as complexity and assertiveness Xu et al. (2023); Hosking et al. (2023). Singhal et al. (2023) study a high correlation between response length and RLHF training. Motivated by these studies, we integrate a set of interpretable features, such as length incentive, repetition penalty, query relevance, and distinction by query type, to understand the behaviors of the learned proxy reward models.
3 Experimental Setup
3.1 Reverse Reward Engineering
Inspired by previous studies on superficial biases in human feedback Singhal et al. (2023); Xu et al. (2023); Hosking et al. (2023), we aim to explore whether there would be a reward function with interpretable features that can maximize the “ground truth” reward signal in reverse, as shown in Figure 2. Started by the length bias Singhal et al. (2023); Chen et al. (2024b), we progressively design the additional features to achieve the goal. More formally, we first perform RL training, i.e., Proximal Policy Optimization (PPO) Schulman et al. (2017), with the designed function as a proxy reward and then consider it a success in reverse engineering if a monotonic relationship between the proxy reward and the ground truth reward is observed in the multiple evaluations over validation set. In other words, if overoptimization Gao et al. (2023) for the proxy reward is detected, it will be considered a failure in reverse engineering.
3.2 Features for White-box Reward
In this subsection, we detail the reward features employed in our experiments on reverse engineering. For a set of pairs, we denote as a query, as a reference response, and as a rollout response of Proximal Policy Optimization (PPO) Schulman et al. (2017); Ziegler et al. (2020).
Length Incentive (LI).
@@ 3.2 Features for White-box Reward
First, we introduce Length Incentive (LI) to exploit the verbosity bias in human preference Singhal et al. (2023); Wang et al. (2024); Chen et al. (2024b). It scores the reward by calculating . Intuitively, this reward function promotes the longer generation regardless of the given query . It is similar to LPPO suggested by Singhal et al. (2023), but we do not normalize the reward with the target length.
Repetition Penalty (RP).
In a preliminary study, solely using LI as a reward function produces undesired repetitions in the generations Welleck et al. (2019). We employ the Repetition penalty (RP) to mitigate unnecessary repetitions in . It checks the unique trigram ratio of , i.e., . Integrating it with the Length Incentive function efficiently prevents redundancy and encourages proper length of outputs, i.e., .
Query Relevance (QR).
We find that combining LI and RP is somewhat effective, but the functions often promote the generation of irrelevant content since both functions are input-agnostic. Hence, we involve the Query Relevance (QR) as one of the reward features. It checks whether the generated response contains off-topic contents for a provided query . We compute the relevance score between the query and sampled response: , where the is pre-trained bi-encoder retriever such as Contriever Izacard et al. (2021).
Reward Branching and Reference Answer Relevance (AR).
@@
The intuition behind the mixture of reward functions is aiming for high relevance between query and generated response while promoting long response length and less repetitions. However, the reward functions, especially LI, could be problematic when the query requires constrained or factual responses. In these cases, consistency of response might be much more important than the long and diverse responses, as discussed in Touvron et al. (2023); Song et al. (2023).
@@
We apply the different relevance rewards according to query type to handle such cases. Specifically, we define two types of queries, Open-ended (OE) and Closed-ended (CE), according to whether they require creative and open-ended or consistent and constrained responses, similar to Song et al. (2023). Then, we include reference answer relevance , where the is the ground-truth reference response of given query. Finally, we denote the final reward design, Reverse Engineered Reward (RER):
, where the indicates the pre-identified query types and the indicates a linear interpolation function mapping the range of AR to the range of for stable RL training 222We empirically find that the multiplication of each reward function works better than the addition of them. Please see Appendix B for more details..
Query-Type Classification.
@@
Deciding whether the given query requires open-ended generation may not be obvious because of its subjectivity in interpreting the range of ‘open-ended’. Nevertheless, we define the meaning of open-ended so that the responses to the query can be relatively anything and creative. For example, “How can I make a good first impression?” could be an open-ended query. On the other side, the closed-ended query indicates the corresponding responses should be consistent and constrained as a closed form. “How many movies are in the Fast and Furious franchise?” would be one of the closed-ended queries. Considering resource constraints, we automatically annotate the query type (OE or CE) used in RER by prompting it to GPT-4 OpenAI (2023) 333We find that about 83% of decisions by GPT-4 are agreed upon when the authors manually validate 100 samples.. Please see Appendix A for more details of the query-type classification.
PPO training.
@@
We use the resulting white-box reward function, RER, to further optimize an SFT model with reinforcement learning. Consequently, the following PPO objective induces the model to maximize RER,
3.3 Evaluation setup
Implementation Details.
We conduct experiments based on two alignment datasets, Anthropic-HH Bai et al. (2022a) and AlpacaFarm Dubois et al. (2023). We perform PPO Ziegler et al. (2020) training with rewards from RER or other reward models at most 5k steps. As in Section 3.2, we employ Contriever Izacard et al. (2021) for scoring relevance scores 444huggingface.co/facebook/contriever. We train SFT model based on LLaMA-2-7B Touvron et al. (2023) with the 161k chosen responses in Anthropic-HH 555huggingface.co/datasets/Anthropic/hh-rlhf. For the PPO training on Anthropic-HH, we select about 23k (train and dev) first-turn queries from ‘Helpful-rejection’ and ‘Helpful-online’ sub-splits for efficient implementation. Also, we employ the pre-trained SFT model for the AlpacaFarm. We use 20k of ‘unlabeled’ split for the PPO training on AlpacaFarm 666huggingface.co/datasets/tatsu-lab/alpaca_farm. More details are in Appendix C.
Gold RM.
As discussed in Gao et al. (2023), it is a non-trivial problem to obtain human labels for the ground-truth reward signal. Hence, following the Gao et al. (2023), we also assume the gold-standard reward model (Gold RM) to evaluate the reward functions. Specifically, we employ one of the strong open-source RMs, StarlingRM-34B 777huggingface.co/Nexusflow/Starling-RM-34B, as the Gold RM. It is trained on the Yi-34B-Chat Young et al. (2024) with Nectar, which is an AI feedback dataset Zhu et al. (2023). This model demonstrates the best performance on various preference evaluation tasks within the recently proposed RewardBench Lambert et al. (2024). We check the reward signal relationship between proxy and gold rewards based on 805 prompts in AlpacaEval Li et al. (2023).
Other Evaluations.
We further evaluate the resulting models on Vicuna-Bench Chiang et al. (2023), AlpacaEval Li et al. (2023), and MT-Bench Zheng et al. (2023). The three benchmarks mainly evaluate the preference of models’ responses leveraging a superior proprietary LLM, GPT-4 OpenAI (2023). In addition, we include Super-NaturalInstructions (SuperNI) Wang et al. (2022) to investigate the alignment tax problem of the resulting models. More evaluation details are in Appendix D.
Relevant Sentence Ratio.
To assess the relevance of the model’s responses for the provided query, we measure the Relevant Sentence Ratio (% Rel. Sent). Specifically, we request GPT-4 OpenAI (2023) to refer to each sentence first and then judge its query relevance as either ‘Relevant’ or ‘Irrelevant’. That is, it indicates the ratio of on-topic sentences in a generated response. The regarding prompt for this evaluation is found in Appendix E.
Baseline RMs.
In addition to the Gold RM, i.e., StarlingRM-34B, used for our analysis, we include three open-source RMs as baselines. OpenAssistant (OASST) is trained with four public feedback datasets, WebGPT Nakano et al. (2021), OpenAI Summary Stiennon et al. (2020), Anthropic-HH Bai et al. (2022a), and SyntheticGPT 888Dahoas/synthetic-instruct-gptj-pairwise, based on DeBERTa-V3-Large He et al. (2021). SteamSHP-XL is another open-source RM trained with SHP Ethayarajh et al. (2023) and Helpfulness subset of the Anthropic-HH. Unlike other feedback datasets, the SHP consists of human-generated responses from an online community, Reddit. It is based on FLAN-T5-XL Chung et al. (2022). Also, we include UltraRM-13B, one of the top open-source RMs trained on UltraFeedback Cui et al. (2023). The UltraFeedback is a 64k AI feedback dataset constructed by GPT-4 OpenAI (2023). Also, it is further fine-tuned with the public feedback datasets, Anthropic-HH, SHP, and OpenAI Summary. We perform PPO training against the reward score by each RM or our reward functions.